1 Dummy

Copyright (c) 1985 Free Software Foundation, Inc; See end for conditions. You are looking at the Emacs tutorial.

Emacs commands generally involve the CONTROL key (sometimes labeled CTRL or CTL) or the META key (sometimes labeled EDIT or ALT). Rather than write that in full each time, we’ll use the following abbreviations:

C-¡chr¿ means hold the CONTROL key while typing the character ¡chr¿ Thus, C-f would be: hold the CONTROL key and type f. M-¡chr¿ means hold the META or EDIT or ALT key down while typing ¡chr¿. If there is no META, EDIT or ALT key, instead press and release the ESC key and then type ¡chr¿. We write ¡ESC¿ for the ESC key.

Important note: to end the Emacs session, type C-x C-c. (Two characters.) The characters ”¿¿” at the left margin indicate directions for you to try using a command. For instance:

[Middle of page left blank for didactic purposes. Text continues below] ¿¿ Now type C-v (View next screen) to move to the next screen. (go

ahead, do it by holding down the CONTROL key while typing v). From now on, you should do this again whenever you finish reading the screen.

Note that there is an overlap of two lines when you move from screen to screen; this provides some continuity so you can continue reading the text.

The first thing that you need to know is how to move around from place to place in the text. You already know how to move forward one screen, with C-v. To move backwards one screen, type M-v (hold down the META key and type v, or type ¡ESC¿v if you do not have a META, EDIT, or ALT key).

¿¿ Try typing M-v and then C-v, a few times.

* SUMMARY ———

The following commands are useful for viewing screenfuls:

C-v Move forward one screenful M-v Move backward one screenful C-l Clear screen and redisplay all the text, moving the text around the cursor to the center of the screen. (That’s CONTROL-L, not CONTROL-1.)

¿¿ Find the cursor, and note what text is near it. Then type C-l. Find the cursor again and notice that the same text is near the cursor now.

* BASIC CURSOR CONTROL ———————-

Moving from screenful to screenful is useful, but how do you move to a specific place within the text on the screen?

There are several ways you can do this. The most basic way is to use the commands C-p, C-b, C-f, and C-n. Each of these commands moves the cursor one row or column in a particular direction on the screen. Here is a table showing these four commands and the directions they move:

Previous line, C-p : : Backward, C-b .... Current cursor position .... Forward, C-f : : Next line, C-n

¿¿ Move the cursor to the line in the middle of that diagram using C-n or C-p. Then type C-l to see the whole diagram centered in the screen.

You’ll probably find it easy to think of these by letter: P for previous, N for next, B for backward and F for forward. These are the basic cursor positioning commands, and you’ll be using them ALL the time, so it would be of great benefit if you learn them now.

¿¿ Do a few C-n’s to bring the cursor down to this line.

¿¿ Move into the line with C-f’s and then up with C-p’s. See what C-p does when the cursor is in the middle of the line.

Each line of text ends with a Newline character, which serves to separate it from the following line. The last line in your file ought to have a Newline at the end (but Emacs does not require it to have one).

¿¿ Try to C-b at the beginning of a line. It should move to the end of the previous line. This is because it moves back across the Newline character.

C-f can move across a Newline just like C-b.

¿¿ Do a few more C-b’s, so you get a feel for where the cursor is. Then do C-f’s to return to the end of the line. Then do one more C-f to move to the following line.

When you move past the top or bottom of the screen, the text beyond the edge shifts onto the screen. This is called ”scrolling”. It enables Emacs to move the cursor to the specified place in the text without moving it off the screen.

¿¿ Try to move the cursor off the bottom of the screen with C-n, and see what happens.

If moving by characters is too slow, you can move by words. M-f (METAf) moves forward a word and M-b moves back a word.

¿¿ Type a few M-f’s and M-b’s.

When you are in the middle of a word, M-f moves to the end of the word. When you are in whitespace between words, M-f moves to the end of the following word. M-b works likewise in the opposite direction.

¿¿ Type M-f and M-b a few times, interspersed with C-f’s and C-b’s so that you can observe the action of M-f and M-b from various places inside and between words.

Notice the parallel between C-f and C-b on the one hand, and M-f and M-b on the other hand. Very often Meta characters are used for operations related to the units defined by language (words, sentences, paragraphs), while Control characters operate on basic units that are independent of what you are editing (characters, lines, etc).

This parallel applies between lines and sentences: C-a and C-e move to the beginning or end of a line, and M-a and M-e move to the beginning or end of a sentence.

¿¿ Try a couple of C-a’s, and then a couple of C-e’s. Try a couple of M-a’s, and then a couple of M-e’s.

See how repeated C-a’s do nothing, but repeated M-a’s keep moving one more sentence. Although these are not quite analogous, each one seems natural.

The location of the cursor in the text is also called ”point”. To paraphrase, the cursor shows on the screen where point is located in the text.

Here is a summary of simple cursor-moving operations, including the word and sentence moving commands:

C-f Move forward a character C-b Move backward a character M-f Move forward a word M-b Move backward a word C-n Move to next line C-p Move to previous line C-a Move to beginning of line C-e Move to end of line M-a Move back to beginning of sentence M-e Move forward to end of

sentence

¿¿ Try all of these commands now a few times for practice. These are the most often used commands.

Two other important cursor motion commands are M-¡ (META Lessthan), which moves to the beginning of the whole text, and M-¿ (META Greater-than), which moves to the end of the whole text.

On most terminals, the ”¡” is above the comma, so you must use the shift key to type it. On these terminals you must use the shift key to type M-¡ also; without the shift key, you would be typing M-comma.

¿¿ Try M-¡ now, to move to the beginning of the tutorial. Then use C-v repeatedly to move back here.

¿¿ Try M-¿ now, to move to the end of the tutorial. Then use M-v repeatedly to move back here.

You can also move the cursor with the arrow keys, if your terminal has arrow keys. We recommend learning C-b, C-f, C-n and C-p for three reasons. First, they work on all kinds of terminals. Second, once you gain practice at using Emacs, you will find that typing these Control characters is faster than typing the arrow keys (because you do not have to move your hands away from touch-typing position). Third, once you form the habit of using these Control character commands, you can easily learn to use other advanced cursor motion commands as well.

Most Emacs commands accept a numeric argument; for most commands, this serves as a repeat-count. The way you give a command a repeat count is by typing C-u and then the digits before you type the command. If you have a META (or EDIT or ALT) key, there is another, alternative way to enter a numeric argument: type the digits while holding down the META key. We recommend learning the C-u method because it works on any terminal. The numeric argument is also called a ”prefix argument”, because you type the argument before the command it applies to.

For instance, C-u 8 C-f moves forward eight characters.

¿¿ Try using C-n or C-p with a numeric argument, to move the cursor to a line near this one with just one command.

Most commands use the numeric argument as a repeat count, but some commands use it in some other way. Several commands (but none of those you have learned so far) use it as a flag–the presence of a prefix argument, regardless of its value, makes the command do something different.

C-v and M-v are another kind of exception. When given an argument, they scroll the screen up or down by that many lines, rather than by a screenful. For example, C-u 8 C-v scrolls the screen by 8 lines.

¿¿ Try typing C-u 8 C-v now.

This should have scrolled the screen up by 8 lines. If you would like to scroll it down again, you can give an argument to M-v.

If you are using a windowed display, such as X11 or MS-Windows, there should be a tall rectangular area called a scroll bar at the left hand side of the Emacs window. You can scroll the text by clicking the mouse in the scroll bar.

¿¿ Try pressing the middle button at the top of the highlighted area within the scroll bar. This should scroll the text to a position determined by how high or low you click.

¿¿ Try moving the mouse up and down, while holding the middle button pressed down. You’ll see that the text scrolls up and down as you move the mouse.

* WHEN EMACS IS HUNG ——————–

If Emacs stops responding to your commands, you can stop it safely by typing C-g. You can use C-g to stop a command which is taking too long to execute.

You can also use C-g to discard a numeric argument or the beginning of a command that you do not want to finish.

¿¿ Type C-u 100 to make a numeric arg of 100, then type C-g. Now type C-f. It should move just one character, because you canceled the argument with C-g.

If you have typed an ¡ESC¿ by mistake, you can get rid of it with a C-g.

* DISABLED COMMANDS ——————-

Some Emacs commands are ”disabled” so that beginning users cannot use them by accident.

If you type one of the disabled commands, Emacs displays a message saying what the command was, and asking you whether you want to go ahead and execute the command.

If you really want to try the command, type Space in answer to the question. Normally, if you do not want to execute the disabled command, answer the question with ”n”.

¿¿ Type C-x C-l (which is a disabled command), then type n to answer the question.

* WINDOWS ———

Emacs can have several windows, each displaying its own text. We will explain later on how to use multiple windows. Right now we want to explain how to get rid of extra windows and go back to basic one-window editing. It is simple:

C-x 1 One window (i.e., kill all other windows).

That is CONTROL-x followed by the digit 1. C-x 1 expands the window which contains the cursor, to occupy the full screen. It deletes all other windows.

¿¿ Move the cursor to this line and type C-u 0 C-l. ¿¿ Type CONTROL-h k CONTROL-f. See how this window shrinks, while a new one appears to display documentation on the CONTROL-f command.

¿¿ Type C-x 1 and see the documentation listing window disappear. This command is unlike the other commands you have learned in that it

consists of two characters. It starts with the character CONTROL-x. There is a whole series of commands that start with CONTROL-x; many of them have to do with windows, files, buffers, and related things. These commands are two, three or four characters long.

* INSERTING AND DELETING ————————

If you want to insert text, just type the text. Characters which you can see, such as A, 7, *, etc. are taken by Emacs as text and inserted immediately. Type ¡Return¿ (the carriage-return key) to insert a Newline character.

You can delete the last character you typed by typing ¡Delete¿. ¡Delete¿ is a key on the keyboard, which may be labeled ”Del”. In some cases, the ”Backspace” key serves as ¡Delete¿, but not always!

More generally, ¡Delete¿ deletes the character immediately before the current cursor position.

¿¿ Do this now–type a few characters, then delete them by typing ¡Delete¿ a few times. Don’t worry about this file being changed; you will not alter the master tutorial. This is your personal copy of it.

When a line of text gets too big for one line on the screen, the line of text is ”continued” onto a second screen line. A backslash (”¨) (or, if you’re using a windowed display, a little curved arrow) at the right margin indicates a line which has been continued.

¿¿ Insert text until you reach the right margin, and keep on inserting. You’ll see a continuation line appear.

¿¿ Use ¡Delete¿s to delete the text until the line fits on one screen line again. The continuation line goes away.

You can delete a Newline character just like any other character. Deleting the Newline character between two lines merges them into one line. If the resulting combined line is too long to fit in the screen width, it will be displayed with a continuation line.

¿¿ Move the cursor to the beginning of a line and type ¡Delete¿. This merges that line with the previous line.

¿¿ Type ¡Return¿ to reinsert the Newline you deleted.

Remember that most Emacs commands can be given a repeat count; this includes text characters. Repeating a text character inserts it several times.

¿¿ Try that now – type C-u 8 * to insert ********.

You’ve now learned the most basic way of typing something in Emacs and correcting errors. You can delete by words or lines as well. Here is a summary of the delete operations:

¡Delete¿ delete the character just before the cursor C-d delete the next character after the cursor

M-¡Delete¿ kill the word immediately before the cursor M-d kill the next word after the cursor

C-k kill from the cursor position to end of line M-k kill to the end of the current sentence

Notice that ¡Delete¿ and C-d vs M-¡Delete¿ and M-d extend the parallel started by C-f and M-f (well, ¡Delete¿ is not really a control character, but let’s not worry about that). C-k and M-k are like C-e and M-e, sort of, in that lines are opposite sentences.

You can also kill any part of the buffer with one uniform method. Move to one end of that part, and type C-@ or C-SPC (either one). (SPC is the Space bar.) Move to the other end of that part, and type C-w. That kills all the text between the two positions.

¿¿ Move the cursor to the Y at the start of the previous paragraph. ¿¿ Type C-SPC. Emacs should display a message ”Mark set” at the bottom of the screen. ¿¿ Move the cursor to the n in ”end”, on the second line of the paragraph. ¿¿ Type C-w. This will kill the text starting from the Y, and ending just before the n.

The difference between ”killing” and ”deleting” is that ”killed” text can be reinserted, whereas ”deleted” things cannot be reinserted. Reinsertion of killed text is called ”yanking”. Generally, the commands that can remove a lot of text kill the text (they set up so that you can yank the text), while the commands that remove just one character, or just blank lines and spaces, do deletion (so you cannot yank that text).

¿¿ Move the cursor to the beginning of a line which is not empty. Then type C-k to kill the text on that line. ¿¿ Type C-k a second time. You’ll see that it kills the Newline which follows that line.

Note that a single C-k kills the contents of the line, and a second C-k kills the line itself, and makes all the other lines move up. C-k treats a numeric argument specially: it kills that many lines AND their contents. This is not mere repetition. C-u 2 C-k kills two lines and their newlines; typing C-k twice would not do that.

Bringing back killed text is called ”yanking”. (Think of it as yanking back, or pulling back, some text that was taken away.) You can yank the killed text either at the same place where it was killed, or at some other place in the buffer, or even in a different file. You can yank the text several times, which makes multiple copies of it.

The command for yanking is C-y. It reinserts the last killed text, at the current cursor position.

¿¿ Try it; type C-y to yank the text back.

If you do several C-k’s in a row, all of the killed text is saved together, so that one C-y will yank all of the lines at once.

¿¿ Do this now, type C-k several times.

Now to retrieve that killed text:

¿¿ Type C-y. Then move the cursor down a few lines and type C-y again. You now see how to copy some text.

What do you do if you have some text you want to yank back, and then you kill something else? C-y would yank the more recent kill. But the previous text is not lost. You can get back to it using the M-y command. After you have done C-y to get the most recent kill, typing M-y replaces that yanked text with the previous kill. Typing M-y again and again brings in earlier and earlier kills. When you have reached the text you are looking for, you do not have to do anything to keep it. Just go on with your editing, leaving the yanked text where it is.

If you M-y enough times, you come back to the starting point (the most recent kill).

¿¿ Kill a line, move around, kill another line. Then do C-y to get back the second killed line. Then do M-y and it will be replaced by the first killed line. Do more M-y’s and see what you get. Keep doing them until the second kill line comes back, and then a few more. If you like, you can try giving M-y positive and negative arguments.

* UNDO ——

If you make a change to the text, and then decide that it was a mistake, you can undo the change with the undo command, C-x u.

Normally, C-x u undoes the changes made by one command; if you repeat the C-x u several times in a row, each repetition undoes one additional command.

But there are two exceptions: commands that do not change the text do not count (this includes cursor motion commands and scrolling command), and self-inserting characters are usually handled in groups of up to 20. (This is to reduce the number of C-x u’s you have to type to undo insertion of text.)

¿¿ Kill this line with C-k, then type C-x u and it should reappear.

C-_{i}sanalternativeundocommand; itworksjustthesameasC−xu, butitiseasiertotypeseveraltimesinarow.T hedisadvantageofC−_{i} sthatonsomekeyboardsitisnotobvioushowtotypeit.T hatiswhyweprovideC−xuaswell.Onsometerminals, youcantypeC−_{b} ytyping/whileholdingdownCONT ROL.

A numeric argument to C-_{o}rC − xuactsasarepeatcount.

You can undo deletion of text just as you can undo killing of text. The distinction between killing something and deleting it affects whether you can yank it with C-y; it makes no difference for undo.

* FILES ——-

In order to make the text you edit permanent, you must put it in a file. Otherwise, it will go away when your invocation of Emacs goes away. In order to put your text in a file, you must ”find” the file before you enter the text. (This is also called ”visiting” the file.)

Finding a file means that you see the contents of the file within Emacs. In many ways, it is as if you were editing the file itself. However, the changes you make using Emacs do not become permanent until you ”save” the file. This is so you can avoid leaving a half-changed file on the system when you do not want to. Even when you save, Emacs leaves the original file under a changed name in case you later decide that your changes were a mistake.

If you look near the bottom of the screen you will see a line that begins and ends with dashes, and starts with ”–:– TUTORIAL” or something like that. This part of the screen normally shows the name of the file that you are visiting. Right now, you are visiting a file called ”TUTORIAL” which is your personal scratch copy of the Emacs tutorial. When you find a file with Emacs, that file’s name will appear in that precise spot.

One special thing about the command for finding a file is that you have to say what file name you want. We say the command ”reads an argument from the terminal” (in this case, the argument is the name of the file). After you type the command

C-x C-f Find a file

Emacs asks you to type the file name. The file name you type appears on the bottom line of the screen. The bottom line is called the minibuffer when it is used for this sort of input. You can use ordinary Emacs editing commands to edit the file name.

While you are entering the file name (or any minibuffer input), you can cancel the command with C-g.

¿¿ Type C-x C-f, then type C-g. This cancels the minibuffer, and also cancels the C-x C-f command that was using the minibuffer. So you do not find any file.

When you have finished entering the file name, type ¡Return¿ to terminate it. Then C-x C-f command goes to work, and finds the file you chose. The minibuffer disappears when the C-x C-f command is finished.

In a little while the file contents appear on the screen, and you can edit the contents. When you wish to make your changes permanent, type the command

C-x C-s Save the file

This copies the text within Emacs into the file. The first time you do this, Emacs renames the original file to a new name so that it is not lost. The new name is made by adding ” ” to the end of the original file’s name.

When saving is finished, Emacs displays the name of the file written. You should save fairly often, so that you will not lose very much work if the system should crash.

¿¿ Type C-x C-s, saving your copy of the tutorial. This should show ”Wrote ...TUTORIAL” at the bottom of the screen.

NOTE: On some systems, typing C-x C-s will freeze the screen and you will see no further output from Emacs. This indicates that an operating system ”feature” called ”flow control” is intercepting the C-s and not letting it get through to Emacs. To unfreeze the screen, type C-q. Then see the section ”Spontaneous Entry to Incremental Search” in the Emacs manual for advice on dealing with this ”feature”.

You can find an existing file, to view it or edit it. You can also find a file which does not already exist. This is the way to create a file with Emacs: find the file, which will start out empty, and then begin inserting the text for the file. When you ask to ”save” the file, Emacs will really create the file with the text that you have inserted. From then on, you can consider yourself to be editing an already existing file.

* BUFFERS ———

If you find a second file with C-x C-f, the first file remains inside Emacs. You can switch back to it by finding it again with C-x C-f. This way you can get quite a number of files inside Emacs.

¿¿ Create a file named ”foo” by typing C-x C-f foo ¡Return¿. Then insert some text, edit it, and save ”foo” by typing C-x C-s. Finally, type C-x C-f TUTORIAL ¡Return¿ to come back to the tutorial.

Emacs stores each file’s text inside an object called a ”buffer”. Finding a file makes a new buffer inside Emacs. To see a list of the buffers that currently exist in your Emacs job, type

C-x C-b List buffers

¿¿ Try C-x C-b now.

See how each buffer has a name, and it may also have a file name for the file whose contents it holds. Some buffers do not correspond to files. For example, the buffer named ”*Buffer List*” does not have any file. It is the buffer which contains the buffer list that was made by C-x C-b. ANY text you see in an Emacs window is always part of some buffer.

¿¿ Type C-x 1 to get rid of the buffer list.

If you make changes to the text of one file, then find another file, this does not save the first file. Its changes remain inside Emacs, in that file’s buffer. The creation or editing of the second file’s buffer has no effect on the first file’s buffer. This is very useful, but it also means that you need a convenient way to save the first file’s buffer. It would be a nuisance to have to switch back to it with C-x C-f in order to save it with C-x C-s. So we have

C-x s Save some buffers

C-x s asks you about each buffer which contains changes that you have not saved. It asks you, for each such buffer, whether to save the buffer.

¿¿ Insert a line of text, then type C-x s. It should ask you whether to save the buffer named TUTORIAL. Answer yes to the question by typing ”y”.

* EXTENDING THE COMMAND SET ————————— There are many, many more Emacs commands than could possibly be

put on all the control and meta characters. Emacs gets around this with the X (eXtend) command. This comes in two flavors:

C-x Character eXtend. Followed by one character. M-x Named command eXtend. Followed by a long name.

These are commands that are generally useful but used less than the commands you have already learned about. You have already seen two of them: the file commands C-x C-f to Find and C-x C-s to Save. Another example is the command to end the Emacs session–this is the command C-x C-c. (Do not worry about losing changes you have made; C-x C-c offers to save each changed file before it kills the Emacs.)

C-z is the command to exit Emacs *temporarily*–so that you can go back to the same Emacs session afterward.

On systems which allow it, C-z ”suspends” Emacs; that is, it returns to the shell but does not destroy the Emacs. In the most common shells, you can resume Emacs with the ‘fg’ command or with ‘

On systems which do not implement suspending, C-z creates a subshell running under Emacs to give you the chance to run other programs and return to Emacs afterward; it does not truly ”exit” from Emacs. In this case, the shell command ‘exit’ is the usual way to get back to Emacs from the subshell.

The time to use C-x C-c is when you are about to log out. It’s also the right thing to use to exit an Emacs invoked under mail handling programs and other miscellaneous utilities, since they may not know how to cope with suspension of Emacs. In ordinary circumstances, though, if you are not about to log out, it is better to suspend Emacs with C-z instead of exiting Emacs.

There are many C-x commands. Here is a list of the ones you have learned:

C-x C-f Find file. C-x C-s Save file. C-x C-b List buffers. C-x C-c Quit Emacs. C-x 1 Delete all but one window. C-x u Undo.

Named eXtended commands are commands which are used even less frequently, or commands which are used only in certain modes. An example is the command replace-string, which globally replaces one string with another. When you type M-x, Emacs prompts you at the bottom of the screen with M-x and you should type the name of the command; in this case, ”replacestring”. Just type ”repl s¡TAB¿” and Emacs will complete the name. (¡TAB¿ is the Tab key, usually found above the CapsLock or Shift key near the left edge of the keyboard.) End the command name with ¡Return¿.

The replace-string command requires two arguments–the string to be replaced, and the string to replace it with. You must end each argument with ¡Return¿.

¿¿ Move the cursor to the blank line two lines below this one. Then type M-x repl s¡Return¿changed¡Return¿altered¡Return¿.

Notice how this line has changed: you’ve replaced the word c-h-a-n-g-e-d with ”altered” wherever it occurred, after the initial position of the cursor.

* AUTO SAVE ———–

When you have made changes in a file, but you have not saved them yet, they could be lost if your computer crashes. To protect you from this, Emacs periodically writes an ”auto save” file for each file that you are editing. The auto save file name has a at the beginning and the end; for example, if your file is named ”hello.c”, its auto save file’s name is ”hello.c”. When you save the file in the normal way, Emacs deletes its auto save file.

If the computer crashes, you can recover your auto-saved editing by finding the file normally (the file you were editing, not the auto save file) and then typing M-x recover file¡Return¿. When it asks for confirmation, type yes¡Return¿ to go ahead and recover the auto-save data.

* ECHO AREA ———–

If Emacs sees that you are typing multicharacter commands slowly, it shows them to you at the bottom of the screen in an area called the ”echo area”. The echo area contains the bottom line of the screen.

* MODE LINE ———–

The line immediately above the echo area is called the ”mode line”. The mode line says something like this:

–:** TUTORIAL (Fundamental)–L670–58

This line gives useful information about the status of Emacs and the text you are editing.

You already know what the filename means–it is the file you have found. -NNthat NN percent of the text is above the top of the screen. If the top of the file is on the screen, it will say –Top– instead of –00–Bot–. If you are looking at text so small that all of it fits on the screen, the mode line says –All–.

The L and digits indicate position in another way: they give the current line number of point.

The stars near the front mean that you have made changes to the text. Right after you visit or save a file, that part of the mode line shows no stars, just dashes.

The part of the mode line inside the parentheses is to tell you what editing modes you are in. The default mode is Fundamental which is what you are using now. It is an example of a ”major mode”.

Emacs has many different major modes. Some of them are meant for editing different languages and/or kinds of text, such as Lisp mode, Text mode, etc. At any time one and only one major mode is active, and its name can always be found in the mode line just where ”Fundamental” is now.

Each major mode makes a few commands behave differently. For example, there are commands for creating comments in a program, and since each programming language has a different idea of what a comment should look like, each major mode has to insert comments differently. Each major mode is the name of an extended command, which is how you can switch to that mode. For example, M-x fundamental-mode is a command to switch to Fundamental mode.

If you are going to be editing English text, such as this file, you should probably use Text Mode. ¿¿ Type M-x text mode¡Return¿.

Don’t worry, none of the Emacs commands you have learned changes in any great way. But you can observe that M-f and M-b now treat apostrophes as part of words. Previously, in Fundamental mode, M-f and M-b treated apostrophes as word-separators.

Major modes usually make subtle changes like that one: most commands do ”the same job” in each major mode, but they work a little bit differently.

To view documentation on your current major mode, type C-h m.

¿¿ Use C-u C-v once or more to bring this line near the top of screen. ¿¿ Type C-h m, to see how Text mode differs from Fundamental mode. ¿¿ Type C-x 1 to remove the documentation from the screen.

Major modes are called major because there are also minor modes. Minor modes are not alternatives to the major modes, just minor modifications of them. Each minor mode can be turned on or off by itself, independent of all other minor modes, and independent of your major mode. So you can use no minor modes, or one minor mode, or any combination of several minor modes.

One minor mode which is very useful, especially for editing English text, is Auto Fill mode. When this mode is on, Emacs breaks the line in between words automatically whenever you insert text and make a line that is too wide.

You can turn Auto Fill mode on by doing M-x auto fill mode¡Return¿. When the mode is on, you can turn it off again by doing M-x auto fill mode¡Return¿. If the mode is off, this command turns it on, and if the mode is on, this command turns it off. We say that the command ”toggles the mode”.

¿¿ Type M-x auto fill mode¡Return¿ now. Then insert a line of ”asdf ” over again until you see it divide into two lines. You must put in spaces between them because Auto Fill breaks lines only at spaces.

The margin is usually set at 70 characters, but you can change it with the C-x f command. You should give the margin setting you want as a numeric argument.

¿¿ Type C-x f with an argument of 20. (C-u 2 0 C-x f). Then type in some text and see Emacs fill lines of 20 characters with it. Then set the margin back to 70 using C-x f again.

If you make changes in the middle of a paragraph, Auto Fill mode does not re-fill it for you. To re-fill the paragraph, type M-q (META-q) with the cursor inside that paragraph.

¿¿ Move the cursor into the previous paragraph and type M-q.

* SEARCHING ———–

Emacs can do searches for strings (these are groups of contiguous characters or words) either forward through the text or backward through it. Searching for a string is a cursor motion command; it moves the cursor to the next place where that string appears.

The Emacs search command is different from the search commands of most editors, in that it is ”incremental”. This means that the search happens while you type in the string to search for.

The command to initiate a search is C-s for forward search, and C-r for reverse search. BUT WAIT! Don’t try them now.

When you type C-s you’ll notice that the string ”I-search” appears as a prompt in the echo area. This tells you that Emacs is in what is called an incremental search waiting for you to type the thing that you want to search for. ¡Return¿ terminates a search.

¿¿ Now type C-s to start a search. SLOWLY, one letter at a time, type the word ’cursor’, pausing after you type each character to notice what happens to the cursor. Now you have searched for ”cursor”, once. ¿¿ Type C-s again, to search for the next occurrence of ”cursor”. ¿¿ Now type ¡Delete¿ four times and see how the cursor moves. ¿¿ Type ¡Return¿ to terminate the search.

Did you see what happened? Emacs, in an incremental search, tries to go to the occurrence of the string that you’ve typed out so far. To go to the next occurrence of ’cursor’ just type C-s again. If no such occurrence exists, Emacs beeps and tells you the search is currently ”failing”. C-g would also terminate the search.

NOTE: On some systems, typing C-s will freeze the screen and you will see no further output from Emacs. This indicates that an operating system ”feature” called ”flow control” is intercepting the C-s and not letting it get through to Emacs. To unfreeze the screen, type C-q. Then see the section ”Spontaneous Entry to Incremental Search” in the Emacs manual for advice on dealing with this ”feature”.

If you are in the middle of an incremental search and type ¡Delete¿, you’ll notice that the last character in the search string is erased and the search backs up to the last place of the search. For instance, suppose you have typed ”c”, to search for the first occurrence of ”c”. Now if you type ”u”, the cursor will move to the first occurrence of ”cu”. Now type ¡Delete¿. This erases the ”u” from the search string, and the cursor moves back to the first occurrence of ”c”.

If you are in the middle of a search and type a control or meta character (with a few exceptions–characters that are special in a search, such as C-s and C-r), the search is terminated.

The C-s starts a search that looks for any occurrence of the search string AFTER the current cursor position. If you want to search for something earlier in the text, type C-r instead. Everything that we have said about C-s also applies to C-r, except that the direction of the search is reversed.

* MULTIPLE WINDOWS ——————

One of the nice features of Emacs is that you can display more than one window on the screen at the same time.

¿¿ Move the cursor to this line and type C-u 0 C-l (that’s CONTROL-L, not CONTROL-1).

¿¿ Now type C-x 2 which splits the screen into two windows. Both windows display this tutorial. The cursor stays in the top window.

¿¿ Type C-M-v to scroll the bottom window. (If you do not have a real META key, type ESC C-v.)

¿¿ Type C-x o (”o” for ”other”) to move the cursor to the bottom window. ¿¿ Use C-v and M-v in the bottom window to scroll it. Keep reading these directions in the top window.

¿¿ Type C-x o again to move the cursor back to the top window. The cursor in the top window is just where it was before.

You can keep using C-x o to switch between the windows. Each window has its own cursor position, but only one window actually shows the cursor. All the ordinary editing commands apply to the window that the cursor is in. We call this the ”selected window”.

The command C-M-v is very useful when you are editing text in one window and using the other window just for reference. You can keep the cursor always in the window where you are editing, and advance through the other window sequentially with C-M-v.

C-M-v is an example of a CONTROL-META character. If you have a real META key, you can type C-M-v by holding down both CONTROL and META while typing v. It does not matter whether CONTROL or META ”comes first,” because both of these keys act by modifying the characters you type.

If you do not have a real META key, and you use ESC instead, the order does matter: you must type ESC followed by CONTROL-v, because CONTROL-ESC v will not work. This is because ESC is a character in its own right, not a modifier key.

¿¿ Type C-x 1 (in the top window) to get rid of the bottom window. (If you had typed C-x 1 in the bottom window, that would get rid of the

top one. Think of this command as ”Keep just one window–the window I am already in.”)

You do not have to display the same buffer in both windows. If you use C-x C-f to find a file in one window, the other window does not change. You can find a file in each window independently.

Here is another way to use two windows to display two different things: ¿¿ Type C-x 4 C-f followed by the name of one of your files. End with

¡Return¿. See the specified file appear in the bottom window. The cursor goes there, too.

¿¿ Type C-x o to go back to the top window, and C-x 1 to delete the bottom window.

* RECURSIVE EDITING LEVELS ————————–

Sometimes you will get into what is called a ”recursive editing level”. This is indicated by square brackets in the mode line, surrounding the parentheses around the major mode name. For example, you might see [(Fundamental)] instead of (Fundamental).

To get out of the recursive editing level, type ESC ESC ESC. That is an all-purpose ”get out” command. You can also use it for eliminating extra windows, and getting out of the minibuffer.

¿¿ Type M-x to get into a minibuffer; then type ESC ESC ESC to get out.

You cannot use C-g to get out of a recursive editing level. This is because C-g is used for canceling commands and arguments WITHIN the recursive editing level.

* GETTING MORE HELP ——————-

In this tutorial we have tried to supply just enough information to get you started using Emacs. There is so much available in Emacs that it would be impossible to explain it all here. However, you may want to learn more about Emacs since it has many other useful features. Emacs provides commands for reading documentation about Emacs commands. These ”help” commands all start with the character CONTROL-h, which is called ”the Help character”.

To use the Help features, type the C-h character, and then a character saying what kind of help you want. If you are REALLY lost, type C-h ? and Emacs will tell you what kinds of help it can give. If you have typed C-h and decide you do not want any help, just type C-g to cancel it.

(Some sites change the meaning of the character C-h. They really should not do this as a blanket measure for all users, so you have grounds to complain to the system administrator. Meanwhile, if C-h does not display a message about help at the bottom of the screen, try typing the F1 key or M-x help ¡Return¿ instead.)

The most basic HELP feature is C-h c. Type C-h, the character c, and a command character or sequence; then Emacs displays a very brief description of the command.

¿¿ Type C-h c C-p. The message should be something like

C-p runs the command previous-line

This tells you the ”name of the function”. Function names are used mainly for customizing and extending Emacs. But since function names are chosen to indicate what the command does, they can serve also as very brief documentation–sufficient to remind you of commands you have already learned.

Multi-character commands such as C-x C-s and (if you have no META or EDIT or ALT key) ¡ESC¿v are also allowed after C-h c.

To get more information about a command, use C-h k instead of C-h c. ¿¿ Type C-h k C-p. This displays the documentation of the function, as well as its name, in

an Emacs window. When you are finished reading the output, type C-x 1 to get rid of the help text. You do not have to do this right away. You can do some editing while referring to the help text, and then type C-x 1.

Here are some other useful C-h options:

C-h f Describe a function. You type in the name of the function.

¿¿ Try typing C-h f previous-line¡Return¿. This displays all the information Emacs has about the function which implements the C-p command.

A similar command C-h v displays the documentation of variables whose values you can set to customize Emacs behavior. You need to type in the name of the variable when Emacs prompts for it.

C-h a Command Apropos. Type in a keyword and Emacs will list all the commands whose names contain that keyword. These commands can all be invoked with META-x. For some commands, Command Apropos will also list a one or two character sequence which runs the same command.

¿¿ Type C-h a file¡Return¿.

This displays in another window a list of all M-x commands with ”file” in their names. You will see character-commands like C-x C-f listed beside the corresponding command names such as find-file.

¿¿ Type C-M-v to scroll the help window. Do this a few times.

¿¿ Type C-x 1 to delete the help window.

C-h i Read On-line Manuals (a.k.a. Info). This command puts you into a special buffer called ‘*info*’ where you can read on-line manuals for the packages installed on your system. Type m emacs ¡Return¿ to read the Emacs manual. If you have never before used Info, type ? and Emacs will take you on a guided tour of Info mode facilities. Once you are through with this tutorial, you should consult the Emacs Info manual as your primary documentation.

* CONCLUSION ————

State  Capitol
Arizona Phoenix
Montana Helena

Table 1: US State Capitols.

Remember, to exit Emacs permanently use C-x C-c. To exit to a shell temporarily, so that you can come back to Emacs afterward, use C-z.

This tutorial is meant to be understandable to all new users, so if you found something unclear, don’t sit and blame yourself - complain!

* COPYING ———

This tutorial descends from a long line of Emacs tutorials starting with the one written by Stuart Cracraft for the original Emacs.

This version of the tutorial, like GNU Emacs, is copyrighted, and comes with permission to distribute copies on certain conditions:

Copyright (c) 1985, 1996 Free Software Foundation

Permission is granted to anyone to make or distribute verbatim copies of this document as received, in any medium, provided that the copyright notice and permission notice are preserved, and that the distributor grants the recipient permission for further redistribution as permitted by this notice.

Permission is granted to distribute modified versions of this document, or of portions of it, under the above conditions, provided also that they carry prominent notices stating who last altered them.

The conditions for copying Emacs itself are more complex, but in the same spirit. Please read the file COPYING and then do give copies of GNU Emacs to your friends. Help stamp out software obstructionism (”ownership”) by using, writing, and sharing free software!

2 With table

Some time ago the printers used to charge extra money for typesetting tables like Table 1.

3 Dummy 2

Copyright (c) 1985 Free Software Foundation, Inc; See end for conditions. You are looking at the Emacs tutorial.

Emacs commands generally involve the CONTROL key (sometimes labeled CTRL or CTL) or the META key (sometimes labeled EDIT or ALT). Rather than write that in full each time, we’ll use the following abbreviations:

C-¡chr¿ means hold the CONTROL key while typing the character ¡chr¿ Thus, C-f would be: hold the CONTROL key and type f. M-¡chr¿ means hold the META or EDIT or ALT key down while typing ¡chr¿. If there is no META, EDIT or ALT key, instead press and release the ESC key and then type ¡chr¿. We write ¡ESC¿ for the ESC key.

Important note: to end the Emacs session, type C-x C-c. (Two characters.) The characters ”¿¿” at the left margin indicate directions for you to try using a command. For instance:

[Middle of page left blank for didactic purposes. Text continues below] ¿¿ Now type C-v (View next screen) to move to the next screen. (go

ahead, do it by holding down the CONTROL key while typing v). From now on, you should do this again whenever you finish reading the screen.

Note that there is an overlap of two lines when you move from screen to screen; this provides some continuity so you can continue reading the text.

The first thing that you need to know is how to move around from place to place in the text. You already know how to move forward one screen, with C-v. To move backwards one screen, type M-v (hold down the META key and type v, or type ¡ESC¿v if you do not have a META, EDIT, or ALT key).

¿¿ Try typing M-v and then C-v, a few times.

* SUMMARY ———

The following commands are useful for viewing screenfuls:

C-v Move forward one screenful M-v Move backward one screenful C-l Clear screen and redisplay all the text, moving the text around the cursor to the center of the screen. (That’s CONTROL-L, not CONTROL-1.)

¿¿ Find the cursor, and note what text is near it. Then type C-l. Find the cursor again and notice that the same text is near the cursor now.

* BASIC CURSOR CONTROL ———————-

Moving from screenful to screenful is useful, but how do you move to a specific place within the text on the screen?

There are several ways you can do this. The most basic way is to use the commands C-p, C-b, C-f, and C-n. Each of these commands moves the cursor one row or column in a particular direction on the screen. Here is a table showing these four commands and the directions they move:

Previous line, C-p : : Backward, C-b .... Current cursor position .... Forward, C-f : : Next line, C-n

¿¿ Move the cursor to the line in the middle of that diagram using C-n or C-p. Then type C-l to see the whole diagram centered in the screen.

You’ll probably find it easy to think of these by letter: P for previous, N for next, B for backward and F for forward. These are the basic cursor positioning commands, and you’ll be using them ALL the time, so it would be of great benefit if you learn them now.

¿¿ Do a few C-n’s to bring the cursor down to this line.

¿¿ Move into the line with C-f’s and then up with C-p’s. See what C-p does when the cursor is in the middle of the line.

Each line of text ends with a Newline character, which serves to separate it from the following line. The last line in your file ought to have a Newline at the end (but Emacs does not require it to have one).

¿¿ Try to C-b at the beginning of a line. It should move to the end of the previous line. This is because it moves back across the Newline character.

C-f can move across a Newline just like C-b.

¿¿ Do a few more C-b’s, so you get a feel for where the cursor is. Then do C-f’s to return to the end of the line. Then do one more C-f to move to the following line.

When you move past the top or bottom of the screen, the text beyond the edge shifts onto the screen. This is called ”scrolling”. It enables Emacs to move the cursor to the specified place in the text without moving it off the screen.

¿¿ Try to move the cursor off the bottom of the screen with C-n, and see what happens.

If moving by characters is too slow, you can move by words. M-f (METAf) moves forward a word and M-b moves back a word.

¿¿ Type a few M-f’s and M-b’s.

When you are in the middle of a word, M-f moves to the end of the word. When you are in whitespace between words, M-f moves to the end of the following word. M-b works likewise in the opposite direction.

¿¿ Type M-f and M-b a few times, interspersed with C-f’s and C-b’s so that you can observe the action of M-f and M-b from various places inside and between words.

Notice the parallel between C-f and C-b on the one hand, and M-f and M-b on the other hand. Very often Meta characters are used for operations related to the units defined by language (words, sentences, paragraphs), while Control characters operate on basic units that are independent of what you are editing (characters, lines, etc).

This parallel applies between lines and sentences: C-a and C-e move to the beginning or end of a line, and M-a and M-e move to the beginning or end of a sentence.

¿¿ Try a couple of C-a’s, and then a couple of C-e’s. Try a couple of M-a’s, and then a couple of M-e’s.

See how repeated C-a’s do nothing, but repeated M-a’s keep moving one more sentence. Although these are not quite analogous, each one seems natural.

The location of the cursor in the text is also called ”point”. To paraphrase, the cursor shows on the screen where point is located in the text.

Here is a summary of simple cursor-moving operations, including the word and sentence moving commands:

C-f Move forward a character C-b Move backward a character M-f Move forward a word M-b Move backward a word C-n Move to next line C-p Move to previous line C-a Move to beginning of line C-e Move to end of line M-a Move back to beginning of sentence M-e Move forward to end of

sentence

¿¿ Try all of these commands now a few times for practice. These are the most often used commands.

Two other important cursor motion commands are M-¡ (META Lessthan), which moves to the beginning of the whole text, and M-¿ (META Greater-than), which moves to the end of the whole text.

On most terminals, the ”¡” is above the comma, so you must use the shift key to type it. On these terminals you must use the shift key to type M-¡ also; without the shift key, you would be typing M-comma.

¿¿ Try M-¡ now, to move to the beginning of the tutorial. Then use C-v repeatedly to move back here.

¿¿ Try M-¿ now, to move to the end of the tutorial. Then use M-v repeatedly to move back here.

You can also move the cursor with the arrow keys, if your terminal has arrow keys. We recommend learning C-b, C-f, C-n and C-p for three reasons. First, they work on all kinds of terminals. Second, once you gain practice at using Emacs, you will find that typing these Control characters is faster than typing the arrow keys (because you do not have to move your hands away from touch-typing position). Third, once you form the habit of using these Control character commands, you can easily learn to use other advanced cursor motion commands as well.

Most Emacs commands accept a numeric argument; for most commands, this serves as a repeat-count. The way you give a command a repeat count is by typing C-u and then the digits before you type the command. If you have a META (or EDIT or ALT) key, there is another, alternative way to enter a numeric argument: type the digits while holding down the META key. We recommend learning the C-u method because it works on any terminal. The numeric argument is also called a ”prefix argument”, because you type the argument before the command it applies to.

For instance, C-u 8 C-f moves forward eight characters.

¿¿ Try using C-n or C-p with a numeric argument, to move the cursor to a line near this one with just one command.

Most commands use the numeric argument as a repeat count, but some commands use it in some other way. Several commands (but none of those you have learned so far) use it as a flag–the presence of a prefix argument, regardless of its value, makes the command do something different.

C-v and M-v are another kind of exception. When given an argument, they scroll the screen up or down by that many lines, rather than by a screenful. For example, C-u 8 C-v scrolls the screen by 8 lines.

¿¿ Try typing C-u 8 C-v now.

This should have scrolled the screen up by 8 lines. If you would like to scroll it down again, you can give an argument to M-v.

If you are using a windowed display, such as X11 or MS-Windows, there should be a tall rectangular area called a scroll bar at the left hand side of the Emacs window. You can scroll the text by clicking the mouse in the scroll bar.

¿¿ Try pressing the middle button at the top of the highlighted area within the scroll bar. This should scroll the text to a position determined by how high or low you click.

¿¿ Try moving the mouse up and down, while holding the middle button pressed down. You’ll see that the text scrolls up and down as you move the mouse.

* WHEN EMACS IS HUNG ——————–

If Emacs stops responding to your commands, you can stop it safely by typing C-g. You can use C-g to stop a command which is taking too long to execute.

You can also use C-g to discard a numeric argument or the beginning of a command that you do not want to finish.

¿¿ Type C-u 100 to make a numeric arg of 100, then type C-g. Now type C-f. It should move just one character, because you canceled the argument with C-g.

If you have typed an ¡ESC¿ by mistake, you can get rid of it with a C-g.

* DISABLED COMMANDS ——————-

Some Emacs commands are ”disabled” so that beginning users cannot use them by accident.

If you type one of the disabled commands, Emacs displays a message saying what the command was, and asking you whether you want to go ahead and execute the command.

If you really want to try the command, type Space in answer to the question. Normally, if you do not want to execute the disabled command, answer the question with ”n”.

¿¿ Type C-x C-l (which is a disabled command), then type n to answer the question.

* WINDOWS ———

Emacs can have several windows, each displaying its own text. We will explain later on how to use multiple windows. Right now we want to explain how to get rid of extra windows and go back to basic one-window editing. It is simple:

C-x 1 One window (i.e., kill all other windows).

That is CONTROL-x followed by the digit 1. C-x 1 expands the window which contains the cursor, to occupy the full screen. It deletes all other windows.

¿¿ Move the cursor to this line and type C-u 0 C-l. ¿¿ Type CONTROL-h k CONTROL-f. See how this window shrinks, while a new one appears to display documentation on the CONTROL-f command.

¿¿ Type C-x 1 and see the documentation listing window disappear. This command is unlike the other commands you have learned in that it

consists of two characters. It starts with the character CONTROL-x. There is a whole series of commands that start with CONTROL-x; many of them have to do with windows, files, buffers, and related things. These commands are two, three or four characters long.

* INSERTING AND DELETING ————————

If you want to insert text, just type the text. Characters which you can see, such as A, 7, *, etc. are taken by Emacs as text and inserted immediately. Type ¡Return¿ (the carriage-return key) to insert a Newline character.

You can delete the last character you typed by typing ¡Delete¿. ¡Delete¿ is a key on the keyboard, which may be labeled ”Del”. In some cases, the ”Backspace” key serves as ¡Delete¿, but not always!

More generally, ¡Delete¿ deletes the character immediately before the current cursor position.

¿¿ Do this now–type a few characters, then delete them by typing ¡Delete¿ a few times. Don’t worry about this file being changed; you will not alter the master tutorial. This is your personal copy of it.

When a line of text gets too big for one line on the screen, the line of text is ”continued” onto a second screen line. A backslash (”¨) (or, if you’re using a windowed display, a little curved arrow) at the right margin indicates a line which has been continued.

¿¿ Insert text until you reach the right margin, and keep on inserting. You’ll see a continuation line appear.

¿¿ Use ¡Delete¿s to delete the text until the line fits on one screen line again. The continuation line goes away.

You can delete a Newline character just like any other character. Deleting the Newline character between two lines merges them into one line. If the resulting combined line is too long to fit in the screen width, it will be displayed with a continuation line.

¿¿ Move the cursor to the beginning of a line and type ¡Delete¿. This merges that line with the previous line.

¿¿ Type ¡Return¿ to reinsert the Newline you deleted.

Remember that most Emacs commands can be given a repeat count; this includes text characters. Repeating a text character inserts it several times.

¿¿ Try that now – type C-u 8 * to insert ********.

You’ve now learned the most basic way of typing something in Emacs and correcting errors. You can delete by words or lines as well. Here is a summary of the delete operations:

¡Delete¿ delete the character just before the cursor C-d delete the next character after the cursor

M-¡Delete¿ kill the word immediately before the cursor M-d kill the next word after the cursor

C-k kill from the cursor position to end of line M-k kill to the end of the current sentence

Notice that ¡Delete¿ and C-d vs M-¡Delete¿ and M-d extend the parallel started by C-f and M-f (well, ¡Delete¿ is not really a control character, but let’s not worry about that). C-k and M-k are like C-e and M-e, sort of, in that lines are opposite sentences.

You can also kill any part of the buffer with one uniform method. Move to one end of that part, and type C-@ or C-SPC (either one). (SPC is the Space bar.) Move to the other end of that part, and type C-w. That kills all the text between the two positions.

¿¿ Move the cursor to the Y at the start of the previous paragraph. ¿¿ Type C-SPC. Emacs should display a message ”Mark set” at the bottom of the screen. ¿¿ Move the cursor to the n in ”end”, on the second line of the paragraph. ¿¿ Type C-w. This will kill the text starting from the Y, and ending just before the n.

The difference between ”killing” and ”deleting” is that ”killed” text can be reinserted, whereas ”deleted” things cannot be reinserted. Reinsertion of killed text is called ”yanking”. Generally, the commands that can remove a lot of text kill the text (they set up so that you can yank the text), while the commands that remove just one character, or just blank lines and spaces, do deletion (so you cannot yank that text).

¿¿ Move the cursor to the beginning of a line which is not empty. Then type C-k to kill the text on that line. ¿¿ Type C-k a second time. You’ll see that it kills the Newline which follows that line.

Note that a single C-k kills the contents of the line, and a second C-k kills the line itself, and makes all the other lines move up. C-k treats a numeric argument specially: it kills that many lines AND their contents. This is not mere repetition. C-u 2 C-k kills two lines and their newlines; typing C-k twice would not do that.

Bringing back killed text is called ”yanking”. (Think of it as yanking back, or pulling back, some text that was taken away.) You can yank the killed text either at the same place where it was killed, or at some other place in the buffer, or even in a different file. You can yank the text several times, which makes multiple copies of it.

The command for yanking is C-y. It reinserts the last killed text, at the current cursor position.

¿¿ Try it; type C-y to yank the text back.

If you do several C-k’s in a row, all of the killed text is saved together, so that one C-y will yank all of the lines at once.

¿¿ Do this now, type C-k several times.

Now to retrieve that killed text:

¿¿ Type C-y. Then move the cursor down a few lines and type C-y again. You now see how to copy some text.

What do you do if you have some text you want to yank back, and then you kill something else? C-y would yank the more recent kill. But the previous text is not lost. You can get back to it using the M-y command. After you have done C-y to get the most recent kill, typing M-y replaces that yanked text with the previous kill. Typing M-y again and again brings in earlier and earlier kills. When you have reached the text you are looking for, you do not have to do anything to keep it. Just go on with your editing, leaving the yanked text where it is.

If you M-y enough times, you come back to the starting point (the most recent kill).

¿¿ Kill a line, move around, kill another line. Then do C-y to get back the second killed line. Then do M-y and it will be replaced by the first killed line. Do more M-y’s and see what you get. Keep doing them until the second kill line comes back, and then a few more. If you like, you can try giving M-y positive and negative arguments.

* UNDO ——

If you make a change to the text, and then decide that it was a mistake, you can undo the change with the undo command, C-x u.

Normally, C-x u undoes the changes made by one command; if you repeat the C-x u several times in a row, each repetition undoes one additional command.

But there are two exceptions: commands that do not change the text do not count (this includes cursor motion commands and scrolling command), and self-inserting characters are usually handled in groups of up to 20. (This is to reduce the number of C-x u’s you have to type to undo insertion of text.)

¿¿ Kill this line with C-k, then type C-x u and it should reappear.

C-_{i}sanalternativeundocommand; itworksjustthesameasC−xu, butitiseasiertotypeseveraltimesinarow.T hedisadvantageofC−_{i} sthatonsomekeyboardsitisnotobvioushowtotypeit.T hatiswhyweprovideC−xuaswell.Onsometerminals, youcantypeC−_{b} ytyping/whileholdingdownCONT ROL.

A numeric argument to C-_{o}rC − xuactsasarepeatcount.

You can undo deletion of text just as you can undo killing of text. The distinction between killing something and deleting it affects whether you can yank it with C-y; it makes no difference for undo.

* FILES ——-

In order to make the text you edit permanent, you must put it in a file. Otherwise, it will go away when your invocation of Emacs goes away. In order to put your text in a file, you must ”find” the file before you enter the text. (This is also called ”visiting” the file.)

Finding a file means that you see the contents of the file within Emacs. In many ways, it is as if you were editing the file itself. However, the changes you make using Emacs do not become permanent until you ”save” the file. This is so you can avoid leaving a half-changed file on the system when you do not want to. Even when you save, Emacs leaves the original file under a changed name in case you later decide that your changes were a mistake.

If you look near the bottom of the screen you will see a line that begins and ends with dashes, and starts with ”–:– TUTORIAL” or something like that. This part of the screen normally shows the name of the file that you are visiting. Right now, you are visiting a file called ”TUTORIAL” which is your personal scratch copy of the Emacs tutorial. When you find a file with Emacs, that file’s name will appear in that precise spot.

One special thing about the command for finding a file is that you have to say what file name you want. We say the command ”reads an argument from the terminal” (in this case, the argument is the name of the file). After you type the command

C-x C-f Find a file

Emacs asks you to type the file name. The file name you type appears on the bottom line of the screen. The bottom line is called the minibuffer when it is used for this sort of input. You can use ordinary Emacs editing commands to edit the file name.

While you are entering the file name (or any minibuffer input), you can cancel the command with C-g.

¿¿ Type C-x C-f, then type C-g. This cancels the minibuffer, and also cancels the C-x C-f command that was using the minibuffer. So you do not find any file.

When you have finished entering the file name, type ¡Return¿ to terminate it. Then C-x C-f command goes to work, and finds the file you chose. The minibuffer disappears when the C-x C-f command is finished.

In a little while the file contents appear on the screen, and you can edit the contents. When you wish to make your changes permanent, type the command

C-x C-s Save the file

This copies the text within Emacs into the file. The first time you do this, Emacs renames the original file to a new name so that it is not lost. The new name is made by adding ” ” to the end of the original file’s name.

When saving is finished, Emacs displays the name of the file written. You should save fairly often, so that you will not lose very much work if the system should crash.

¿¿ Type C-x C-s, saving your copy of the tutorial. This should show ”Wrote ...TUTORIAL” at the bottom of the screen.

NOTE: On some systems, typing C-x C-s will freeze the screen and you will see no further output from Emacs. This indicates that an operating system ”feature” called ”flow control” is intercepting the C-s and not letting it get through to Emacs. To unfreeze the screen, type C-q. Then see the section ”Spontaneous Entry to Incremental Search” in the Emacs manual for advice on dealing with this ”feature”.

You can find an existing file, to view it or edit it. You can also find a file which does not already exist. This is the way to create a file with Emacs: find the file, which will start out empty, and then begin inserting the text for the file. When you ask to ”save” the file, Emacs will really create the file with the text that you have inserted. From then on, you can consider yourself to be editing an already existing file.

* BUFFERS ———

If you find a second file with C-x C-f, the first file remains inside Emacs. You can switch back to it by finding it again with C-x C-f. This way you can get quite a number of files inside Emacs.

¿¿ Create a file named ”foo” by typing C-x C-f foo ¡Return¿. Then insert some text, edit it, and save ”foo” by typing C-x C-s. Finally, type C-x C-f TUTORIAL ¡Return¿ to come back to the tutorial.

Emacs stores each file’s text inside an object called a ”buffer”. Finding a file makes a new buffer inside Emacs. To see a list of the buffers that currently exist in your Emacs job, type

C-x C-b List buffers

¿¿ Try C-x C-b now.

See how each buffer has a name, and it may also have a file name for the file whose contents it holds. Some buffers do not correspond to files. For example, the buffer named ”*Buffer List*” does not have any file. It is the buffer which contains the buffer list that was made by C-x C-b. ANY text you see in an Emacs window is always part of some buffer.

¿¿ Type C-x 1 to get rid of the buffer list.

If you make changes to the text of one file, then find another file, this does not save the first file. Its changes remain inside Emacs, in that file’s buffer. The creation or editing of the second file’s buffer has no effect on the first file’s buffer. This is very useful, but it also means that you need a convenient way to save the first file’s buffer. It would be a nuisance to have to switch back to it with C-x C-f in order to save it with C-x C-s. So we have

C-x s Save some buffers

C-x s asks you about each buffer which contains changes that you have not saved. It asks you, for each such buffer, whether to save the buffer.

¿¿ Insert a line of text, then type C-x s. It should ask you whether to save the buffer named TUTORIAL. Answer yes to the question by typing ”y”.

* EXTENDING THE COMMAND SET ————————— There are many, many more Emacs commands than could possibly be

put on all the control and meta characters. Emacs gets around this with the X (eXtend) command. This comes in two flavors:

C-x Character eXtend. Followed by one character. M-x Named command eXtend. Followed by a long name.

These are commands that are generally useful but used less than the commands you have already learned about. You have already seen two of them: the file commands C-x C-f to Find and C-x C-s to Save. Another example is the command to end the Emacs session–this is the command C-x C-c. (Do not worry about losing changes you have made; C-x C-c offers to save each changed file before it kills the Emacs.)

C-z is the command to exit Emacs *temporarily*–so that you can go back to the same Emacs session afterward.

On systems which allow it, C-z ”suspends” Emacs; that is, it returns to the shell but does not destroy the Emacs. In the most common shells, you can resume Emacs with the ‘fg’ command or with ‘

On systems which do not implement suspending, C-z creates a subshell running under Emacs to give you the chance to run other programs and return to Emacs afterward; it does not truly ”exit” from Emacs. In this case, the shell command ‘exit’ is the usual way to get back to Emacs from the subshell.

The time to use C-x C-c is when you are about to log out. It’s also the right thing to use to exit an Emacs invoked under mail handling programs and other miscellaneous utilities, since they may not know how to cope with suspension of Emacs. In ordinary circumstances, though, if you are not about to log out, it is better to suspend Emacs with C-z instead of exiting Emacs.

There are many C-x commands. Here is a list of the ones you have learned:

C-x C-f Find file. C-x C-s Save file. C-x C-b List buffers. C-x C-c Quit Emacs. C-x 1 Delete all but one window. C-x u Undo.

Named eXtended commands are commands which are used even less frequently, or commands which are used only in certain modes. An example is the command replace-string, which globally replaces one string with another. When you type M-x, Emacs prompts you at the bottom of the screen with M-x and you should type the name of the command; in this case, ”replacestring”. Just type ”repl s¡TAB¿” and Emacs will complete the name. (¡TAB¿ is the Tab key, usually found above the CapsLock or Shift key near the left edge of the keyboard.) End the command name with ¡Return¿.

The replace-string command requires two arguments–the string to be replaced, and the string to replace it with. You must end each argument with ¡Return¿.

¿¿ Move the cursor to the blank line two lines below this one. Then type M-x repl s¡Return¿changed¡Return¿altered¡Return¿.

Notice how this line has changed: you’ve replaced the word c-h-a-n-g-e-d with ”altered” wherever it occurred, after the initial position of the cursor.

* AUTO SAVE ———–

When you have made changes in a file, but you have not saved them yet, they could be lost if your computer crashes. To protect you from this, Emacs periodically writes an ”auto save” file for each file that you are editing. The auto save file name has a at the beginning and the end; for example, if your file is named ”hello.c”, its auto save file’s name is ”hello.c”. When you save the file in the normal way, Emacs deletes its auto save file.

If the computer crashes, you can recover your auto-saved editing by finding the file normally (the file you were editing, not the auto save file) and then typing M-x recover file¡Return¿. When it asks for confirmation, type yes¡Return¿ to go ahead and recover the auto-save data.

* ECHO AREA ———–

If Emacs sees that you are typing multicharacter commands slowly, it shows them to you at the bottom of the screen in an area called the ”echo area”. The echo area contains the bottom line of the screen.

* MODE LINE ———–

The line immediately above the echo area is called the ”mode line”. The mode line says something like this:

–:** TUTORIAL (Fundamental)–L670–58

This line gives useful information about the status of Emacs and the text you are editing.

You already know what the filename means–it is the file you have found. -NNthat NN percent of the text is above the top of the screen. If the top of the file is on the screen, it will say –Top– instead of –00–Bot–. If you are looking at text so small that all of it fits on the screen, the mode line says –All–.

The L and digits indicate position in another way: they give the current line number of point.

The stars near the front mean that you have made changes to the text. Right after you visit or save a file, that part of the mode line shows no stars, just dashes.

The part of the mode line inside the parentheses is to tell you what editing modes you are in. The default mode is Fundamental which is what you are using now. It is an example of a ”major mode”.

Emacs has many different major modes. Some of them are meant for editing different languages and/or kinds of text, such as Lisp mode, Text mode, etc. At any time one and only one major mode is active, and its name can always be found in the mode line just where ”Fundamental” is now.

Each major mode makes a few commands behave differently. For example, there are commands for creating comments in a program, and since each programming language has a different idea of what a comment should look like, each major mode has to insert comments differently. Each major mode is the name of an extended command, which is how you can switch to that mode. For example, M-x fundamental-mode is a command to switch to Fundamental mode.

If you are going to be editing English text, such as this file, you should probably use Text Mode. ¿¿ Type M-x text mode¡Return¿.

Don’t worry, none of the Emacs commands you have learned changes in any great way. But you can observe that M-f and M-b now treat apostrophes as part of words. Previously, in Fundamental mode, M-f and M-b treated apostrophes as word-separators.

Major modes usually make subtle changes like that one: most commands do ”the same job” in each major mode, but they work a little bit differently.

To view documentation on your current major mode, type C-h m.

¿¿ Use C-u C-v once or more to bring this line near the top of screen. ¿¿ Type C-h m, to see how Text mode differs from Fundamental mode. ¿¿ Type C-x 1 to remove the documentation from the screen.

Major modes are called major because there are also minor modes. Minor modes are not alternatives to the major modes, just minor modifications of them. Each minor mode can be turned on or off by itself, independent of all other minor modes, and independent of your major mode. So you can use no minor modes, or one minor mode, or any combination of several minor modes.

One minor mode which is very useful, especially for editing English text, is Auto Fill mode. When this mode is on, Emacs breaks the line in between words automatically whenever you insert text and make a line that is too wide.

You can turn Auto Fill mode on by doing M-x auto fill mode¡Return¿. When the mode is on, you can turn it off again by doing M-x auto fill mode¡Return¿. If the mode is off, this command turns it on, and if the mode is on, this command turns it off. We say that the command ”toggles the mode”.

¿¿ Type M-x auto fill mode¡Return¿ now. Then insert a line of ”asdf ” over again until you see it divide into two lines. You must put in spaces between them because Auto Fill breaks lines only at spaces.

The margin is usually set at 70 characters, but you can change it with the C-x f command. You should give the margin setting you want as a numeric argument.

¿¿ Type C-x f with an argument of 20. (C-u 2 0 C-x f). Then type in some text and see Emacs fill lines of 20 characters with it. Then set the margin back to 70 using C-x f again.

If you make changes in the middle of a paragraph, Auto Fill mode does not re-fill it for you. To re-fill the paragraph, type M-q (META-q) with the cursor inside that paragraph.

¿¿ Move the cursor into the previous paragraph and type M-q.

* SEARCHING ———–

Emacs can do searches for strings (these are groups of contiguous characters or words) either forward through the text or backward through it. Searching for a string is a cursor motion command; it moves the cursor to the next place where that string appears.

The Emacs search command is different from the search commands of most editors, in that it is ”incremental”. This means that the search happens while you type in the string to search for.

The command to initiate a search is C-s for forward search, and C-r for reverse search. BUT WAIT! Don’t try them now.

When you type C-s you’ll notice that the string ”I-search” appears as a prompt in the echo area. This tells you that Emacs is in what is called an incremental search waiting for you to type the thing that you want to search for. ¡Return¿ terminates a search.

¿¿ Now type C-s to start a search. SLOWLY, one letter at a time, type the word ’cursor’, pausing after you type each character to notice what happens to the cursor. Now you have searched for ”cursor”, once. ¿¿ Type C-s again, to search for the next occurrence of ”cursor”. ¿¿ Now type ¡Delete¿ four times and see how the cursor moves. ¿¿ Type ¡Return¿ to terminate the search.

Did you see what happened? Emacs, in an incremental search, tries to go to the occurrence of the string that you’ve typed out so far. To go to the next occurrence of ’cursor’ just type C-s again. If no such occurrence exists, Emacs beeps and tells you the search is currently ”failing”. C-g would also terminate the search.

NOTE: On some systems, typing C-s will freeze the screen and you will see no further output from Emacs. This indicates that an operating system ”feature” called ”flow control” is intercepting the C-s and not letting it get through to Emacs. To unfreeze the screen, type C-q. Then see the section ”Spontaneous Entry to Incremental Search” in the Emacs manual for advice on dealing with this ”feature”.

If you are in the middle of an incremental search and type ¡Delete¿, you’ll notice that the last character in the search string is erased and the search backs up to the last place of the search. For instance, suppose you have typed ”c”, to search for the first occurrence of ”c”. Now if you type ”u”, the cursor will move to the first occurrence of ”cu”. Now type ¡Delete¿. This erases the ”u” from the search string, and the cursor moves back to the first occurrence of ”c”.

If you are in the middle of a search and type a control or meta character (with a few exceptions–characters that are special in a search, such as C-s and C-r), the search is terminated.

The C-s starts a search that looks for any occurrence of the search string AFTER the current cursor position. If you want to search for something earlier in the text, type C-r instead. Everything that we have said about C-s also applies to C-r, except that the direction of the search is reversed.

* MULTIPLE WINDOWS ——————

One of the nice features of Emacs is that you can display more than one window on the screen at the same time.

¿¿ Move the cursor to this line and type C-u 0 C-l (that’s CONTROL-L, not CONTROL-1).

¿¿ Now type C-x 2 which splits the screen into two windows. Both windows display this tutorial. The cursor stays in the top window.

¿¿ Type C-M-v to scroll the bottom window. (If you do not have a real META key, type ESC C-v.)

¿¿ Type C-x o (”o” for ”other”) to move the cursor to the bottom window. ¿¿ Use C-v and M-v in the bottom window to scroll it. Keep reading these directions in the top window.

¿¿ Type C-x o again to move the cursor back to the top window. The cursor in the top window is just where it was before.

You can keep using C-x o to switch between the windows. Each window has its own cursor position, but only one window actually shows the cursor. All the ordinary editing commands apply to the window that the cursor is in. We call this the ”selected window”.

The command C-M-v is very useful when you are editing text in one window and using the other window just for reference. You can keep the cursor always in the window where you are editing, and advance through the other window sequentially with C-M-v.

C-M-v is an example of a CONTROL-META character. If you have a real META key, you can type C-M-v by holding down both CONTROL and META while typing v. It does not matter whether CONTROL or META ”comes first,” because both of these keys act by modifying the characters you type.

If you do not have a real META key, and you use ESC instead, the order does matter: you must type ESC followed by CONTROL-v, because CONTROL-ESC v will not work. This is because ESC is a character in its own right, not a modifier key.

¿¿ Type C-x 1 (in the top window) to get rid of the bottom window. (If you had typed C-x 1 in the bottom window, that would get rid of the

top one. Think of this command as ”Keep just one window–the window I am already in.”)

You do not have to display the same buffer in both windows. If you use C-x C-f to find a file in one window, the other window does not change. You can find a file in each window independently.

Here is another way to use two windows to display two different things: ¿¿ Type C-x 4 C-f followed by the name of one of your files. End with

¡Return¿. See the specified file appear in the bottom window. The cursor goes there, too.

¿¿ Type C-x o to go back to the top window, and C-x 1 to delete the bottom window.

* RECURSIVE EDITING LEVELS ————————–

Sometimes you will get into what is called a ”recursive editing level”. This is indicated by square brackets in the mode line, surrounding the parentheses around the major mode name. For example, you might see [(Fundamental)] instead of (Fundamental).

To get out of the recursive editing level, type ESC ESC ESC. That is an all-purpose ”get out” command. You can also use it for eliminating extra windows, and getting out of the minibuffer.

¿¿ Type M-x to get into a minibuffer; then type ESC ESC ESC to get out.

You cannot use C-g to get out of a recursive editing level. This is because C-g is used for canceling commands and arguments WITHIN the recursive editing level.

* GETTING MORE HELP ——————-

In this tutorial we have tried to supply just enough information to get you started using Emacs. There is so much available in Emacs that it would be impossible to explain it all here. However, you may want to learn more about Emacs since it has many other useful features. Emacs provides commands for reading documentation about Emacs commands. These ”help” commands all start with the character CONTROL-h, which is called ”the Help character”.

To use the Help features, type the C-h character, and then a character saying what kind of help you want. If you are REALLY lost, type C-h ? and Emacs will tell you what kinds of help it can give. If you have typed C-h and decide you do not want any help, just type C-g to cancel it.

(Some sites change the meaning of the character C-h. They really should not do this as a blanket measure for all users, so you have grounds to complain to the system administrator. Meanwhile, if C-h does not display a message about help at the bottom of the screen, try typing the F1 key or M-x help ¡Return¿ instead.)

The most basic HELP feature is C-h c. Type C-h, the character c, and a command character or sequence; then Emacs displays a very brief description of the command.

¿¿ Type C-h c C-p. The message should be something like

C-p runs the command previous-line

This tells you the ”name of the function”. Function names are used mainly for customizing and extending Emacs. But since function names are chosen to indicate what the command does, they can serve also as very brief documentation–sufficient to remind you of commands you have already learned.

Multi-character commands such as C-x C-s and (if you have no META or EDIT or ALT key) ¡ESC¿v are also allowed after C-h c.

To get more information about a command, use C-h k instead of C-h c. ¿¿ Type C-h k C-p. This displays the documentation of the function, as well as its name, in

an Emacs window. When you are finished reading the output, type C-x 1 to get rid of the help text. You do not have to do this right away. You can do some editing while referring to the help text, and then type C-x 1.

Here are some other useful C-h options:

C-h f Describe a function. You type in the name of the function.

¿¿ Try typing C-h f previous-line¡Return¿. This displays all the information Emacs has about the function which implements the C-p command.

A similar command C-h v displays the documentation of variables whose values you can set to customize Emacs behavior. You need to type in the name of the variable when Emacs prompts for it.

C-h a Command Apropos. Type in a keyword and Emacs will list all the commands whose names contain that keyword. These commands can all be invoked with META-x. For some commands, Command Apropos will also list a one or two character sequence which runs the same command.

¿¿ Type C-h a file¡Return¿.

This displays in another window a list of all M-x commands with ”file” in their names. You will see character-commands like C-x C-f listed beside the corresponding command names such as find-file.

¿¿ Type C-M-v to scroll the help window. Do this a few times.

¿¿ Type C-x 1 to delete the help window.

C-h i Read On-line Manuals (a.k.a. Info). This command puts you into a special buffer called ‘*info*’ where you can read on-line manuals for the packages installed on your system. Type m emacs ¡Return¿ to read the Emacs manual. If you have never before used Info, type ? and Emacs will take you on a guided tour of Info mode facilities. Once you are through with this tutorial, you should consult the Emacs Info manual as your primary documentation.

* CONCLUSION ————

Remember, to exit Emacs permanently use C-x C-c. To exit to a shell temporarily, so that you can come back to Emacs afterward, use C-z.

This tutorial is meant to be understandable to all new users, so if you found something unclear, don’t sit and blame yourself - complain!

* COPYING ———

This tutorial descends from a long line of Emacs tutorials starting with the one written by Stuart Cracraft for the original Emacs.

This version of the tutorial, like GNU Emacs, is copyrighted, and comes with permission to distribute copies on certain conditions:

Copyright (c) 1985, 1996 Free Software Foundation

Permission is granted to anyone to make or distribute verbatim copies of this document as received, in any medium, provided that the copyright notice and permission notice are preserved, and that the distributor grants the recipient permission for further redistribution as permitted by this notice.

Permission is granted to distribute modified versions of this document, or of portions of it, under the above conditions, provided also that they carry prominent notices stating who last altered them.

The conditions for copying Emacs itself are more complex, but in the same spirit. Please read the file COPYING and then do give copies of GNU Emacs to your friends. Help stamp out software obstructionism (”ownership”) by using, writing, and sharing free software!